OpenStudy (anonymous):

What is the radius of a circle with a circumference of fourteen point one three meters? Use three point one four for pi and show your work.

OpenStudy (anonymous):

im think its 2.25

OpenStudy (austinl):

Circumference is, \(C=2\pi r^2\) \(C=14.13m\) \(\pi=3.14\) So we have, \(14.13=2(3.14)(r^2)\) Solve for r. \(14.13=6.28r^2\) \(2.25=r^2\) \(r=\sqrt{2.25}=1.5\)

OpenStudy (anonymous):

so its 1.5?

OpenStudy (austinl):

It is indeed. You were almost there, you just needed to take the square root to get r, not r^2.

OpenStudy (anonymous):

oh but im trying to fin the radius why od i do square root

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (austinl):

Because the formula for circumference is, \(C=2\pi\color{red}{r^2}\) Radius is \(r\) not \(r^2\). So to get rid of the square, you take the root of it.

OpenStudy (anonymous):

the formula is 2*3.14*r

OpenStudy (austinl):

Oh, oops it is indeed. In that case, it is \(\color{blue}{2.25}\)

OpenStudy (anonymous):

so i was right right?

OpenStudy (anonymous):

anyways can you show me how to work that

OpenStudy (austinl):

\(C=2\pi r\) \(C=14.13m\\ π=3.14\) \(14.13=2(3.14)r~~~\color{green}{\text{distribute}}\\ 14.13=6.28r~~~\color{green}{\text{divide by 6.28}}\\ r=2.25\)

OpenStudy (anonymous):

alrite thx